[Brahms-l] Shift report 20040210 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Tue Feb 10 2004 - 23:49:18 EST
------------------------------------------------------------
Date: Tue Feb 10 23:49:17 2004, Shift: 16:00-24:00
Supervisor: Ionut ARSENE, with: Bjorn SAMSET
Log book#15 pages 52 - 52

Shift Summary:
------------------------------------------------------------
1. Good beam for 4 hours.

2. Data runs: 10254-10257
 The settings for the entire shift:
  FS at 4 degrees, full field.
  MRS at 45 degrees, A2255
 We achieved ~420k trig2 and ~450k trig3 events.
   Pedestals: 10258(tpc pedestals), 10259(non-tpc pedestals)

3. No detector or magnet problems. T5 tripped one time, but when I checked on BRAHVO everything was OK. 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Feb 10 23:50:08 2004

This archive was generated by hypermail 2.1.8 : Tue Feb 10 2004 - 23:50:48 EST